If the torque required to loosen a nut that holds a wheel on a car has a magnitude of 58 N*m, what force

must be exerted at the end of a 35 cm lug wrench to loosen the nut when the angle of the wrench is 560?



1
Expert's answer
2022-01-18T09:35:38-0500

Write the expression:


"T=Fd\sin\theta→F=T/(d\sin\theta)=200\text{ N}."